(MD order Benadryl 25 mg IV every 3 hrs prn for nausea)

Calculate how many mL of Benadryl the patient will receive IV. supply 10mg/mL
Answer #1

Single dose received by patient in 2.5 ml

Benadryl 10 mg in = 1 ml

Benedryl 1 mg in = 1/10 ml

Benedryl 25 mg in = 25/10 = 2.5 ml

Every 3 hours ,so total dose in day = 8 Doses

So , total ml of supplying required = 8*2.5  

20ml benedryl solution supply